OpenStudy (hba):

Find all the local maxima, local minima and saddle points of the function. f(x,y) = x e^y

OpenStudy (hba):

@ganeshie8

OpenStudy (hba):

fx(x,y)=e^y fy(x,y)=xe^y e^y=0 x.e^y=0 y is undefined x=0 How do i continue?

OpenStudy (anonymous):

fx(x,y) = e^y fx(1,0) = e^0 =1 fy(x,y) = x e^y fy(1,0) = (1)e^0 =1 f(1,0) = 1

OpenStudy (anonymous):

f(x,y) = 1 + 1(x-1) + 1 (y-0) = 1+(x-1)+y = x+y --------------------------------------...

OpenStudy (hba):

Where did 1,0 come from?

OpenStudy (anonymous):

Second-degree Taylor polynomial Follow the example given in the link. I have computed the derivative part. Second degree Taylor polynomials pf f(x,y) at the point (1,0)

OpenStudy (anonymous):

∂f/∂x = e^y ∂f/∂y = x e^y ∂^2f/∂x^2 = 0 ∂^2f/∂y^2 = x e^y ∂^2f/∂x∂y = e^y

OpenStudy (phi):

test for a saddle point is fxx * fyy - (fxy)^2 < 0

OpenStudy (hba):

I know that phi but as y is undefined how am i supposed to continue with this?

OpenStudy (anonymous):

exactly .

OpenStudy (hba):

fxx(x,y)=0,fxy(x,y)=xe^y,fyy(x,y)=e^y Iv'e also calculated those things.

OpenStudy (mathmale):

I'd suggest you go back and quickly review the CRITERIA to be used in determining whether or not you have a saddle point or a maximum or minimum. Would you please defend y our statement, "y is undefined x=0" ?

OpenStudy (hba):

e^y=0 Apply ln to both sides ln(e^y) = ln(0) y = undefined since ln(0) is undefined

OpenStudy (hba):

x.e^y=0 x=0

OpenStudy (abb0t):

You don't need my help. You have a team of mathletes helping you here @hba ...

OpenStudy (mathmale):

Let's re-visit the original f(x,y) = x e^y (or f(x,y) = x*e^y ): \[f _{x}=e^y;~f _{xx}=0;~f _{y}=x*e^y;~ f _{yy}=x*e^y\]

OpenStudy (mathmale):

Please look those partial derivatives over and ask yourself whether you agree or disagree. I don't see that you need to apply the 'ln' operator here.

OpenStudy (hba):

I agree with you 100 % @mathmale but we have to put fx=0 and fy=0 and get (x,y) you see

OpenStudy (hba):

So at that point we can determine local maxima, local minima and saddle points using this: fxx * fyy - (fxy)^2 < 0

OpenStudy (hba):

Forget the <0 thing

OpenStudy (phi):

you have a degenerate case, and it looks like there is no max,min or saddle. (though I don't know how to prove this)

OpenStudy (abb0t):

Ok, let me see, if I remember correctly, if you have (a,b), which is a critical point of f(x,y), and the SECOND ORDER PARTIAL DERIV are continuous on the given region, then: D=D(a,b)=f\(\sf_{xx}\)(a,b)f\(\sf_{yy}\)(a,b)-[f\(\sf_{xy}\)(a,b)]\(^2\) If D>0 and f\(\sf_{xx}\) > 0, its relative min @ (a,b) if D>0 and \(\sf_{xx}\) < 0, there theres a relative max @ (a,b) if D<0 then the point (a,b) is a saddle point D = -, then try something else

OpenStudy (hba):

D=0 then test is conclusive

OpenStudy (abb0t):

I'm sure you remember from Calc 1 how to find ur critical points.

OpenStudy (abb0t):

Yes, sorry, D= 0*

OpenStudy (hba):

@abb0t If y wasn't undefined i would have done it

OpenStudy (abb0t):

@amistre64

OpenStudy (shadowlegendx):

Dis calculus doe